# Uni-AIMS: AI-Powered Microscopy Image Analysis

### v1

@@ 4.3.2 Scale Bar and Text Detection
We adopted the YOLOv8m model ([yolov8_ultralytics]) for the joint detection of scale bars and associated text regions. During training, moderate data augmentation techniques such as horizontal flipping, color jittering, and small-angle rotations were applied. The input image size was set to 1280 pixels, and a 1.5$\times$ down-sampling of the image width was performed to accommodate the typical aspect ratios of scale bars.
@@ 4.3.4 Text Recognition and Scale Bar Information Matching
The recognized text content was parsed using regular expressions (pattern matching) to extract effective measurement values and units (e.g., “5 $\mu$m”, “2 nm”). Based on spatial proximity, each recognized text segment was matched to its nearest scale bar.
@@ 4.3.5 Evaluation
Figure 7 shows the error between the scale bar lengths recognized by our model and the ground truth values across different images. Each point represents the error for an individual image, calculated as (recognized length - ground truth length)/ground truth length. The text box in the lower right corner shows key statistics of the absolute errors, including mean and percentile values. The results indicate that the majority of errors are very small, with the mean absolute error being 0.3%. And more than 50% of the data have an identification length that is exactly correct. Even at higher percentiles (75%, 90%, 95%), the errors remain relatively small, indicating consistent performance across most images.
@@ 5.3 User Interaction and Reporting
Figure 8(c) demonstrates the interactive capabilities of the platform. Users can filter particles by specifying a diameter range (e.g., 5–15 $\mu$m) and choosing to exclude incomplete particles located at the image edges. After filtering, only particles meeting the criteria are displayed in the central region of the image. These filtered data are then used for subsequent statistical analyses.

### v2

@@ 4.3.2 Scale Bar and Text Detection
[…]
@@ 4.3.4 Text Recognition and Scale Bar Information Matching
[…]
@@ 4.3.5 Evaluation
Figure 7 shows the error between the scale bar lengths recognized by our model and the ground truth values across different images. Each point represents the error for an individual image, calculated as (recognized length - ground truth length)/ground truth length. The text box in the lower right corner shows key statistics of the absolute errors, including mean and percentile values. The results indicate that the majority of errors are very small, with a mean absolute error of 0.3%. And more than 50% of the data have an identification length that is exactly correct. Even at higher percentiles (75%, 90%, 95%), the errors remain relatively small, indicating consistent performance across most images.
@@ 5.3 User Interaction and Reporting
[…]